# The Sunburst Arc with JWST. IV. The importance of interaction, turbulence, and feedback for Lyman-continuum escape

### v2

@@ 5.1.1 The origin of a broad emission component in H and [O III]
It has long been predicted that stellar winds and supernova feedback would play an important role in clearing the escape paths for Lyman-continuum, both in theoretical works (M. Trebitsch et al., 2017; J. Rosdahl et al., 2018; K. Kakiichi & M. Gronke, 2021) and in observations (T. M. Heckman et al., 2011; J. Chisholm et al., 2017; K. Kim et al., 2020). And R. O. Amorín et al. (2024) have found a strong preponderance of broad in LyC-leaking Green Pea galaxies, although it has been difficult to establish any direct correlations between broad emission components and LyC escape fraction.
@@
R. Mainali et al. (2022) studied the rest-UV and -optical emission in the Sunburst Arc from the ground using the Magellan/MagE and Magellan/FIRE spectrographs, respectively. These authors did find a strong broad component in the emission features of H and [O III] in the LCE spectra, while spectra from non-leaking regions showed a considerably fainter broad feature. However, the broad feature in R. Mainali et al. has FWHM km s-1 or km s-1 somewhere in between the broad- and medium-width features of (R. O. Amorín et al., 2024). Compared to the ground-based spatially unresolved spectra of R. Mainali et al. (2022), the current NIRSpec IFU observations enable spectra to be extracted from spatial regions much more selectively, with lower contamination from nearby regions.
@@
T. E. Rivera-Thorsen et al. (2024); B. Welch et al. (2025) extracted and stacked spectra from multiple lensed images of the LCE, allowing for unprecedented signal/noise ratio and absence of contamination from surrounding regions. Both these works modeled strong, rest-frame optical nebular emission lines using three Gaussian components, yielding FWHM values of km/s; consistent with each other and broadly consistent with typical values of the narrow, medium and broad components found by R. O. Amorín et al. (2024).
@@
In this work, we have only used one component to model the emission lines in each spaxel, so our measured velocities are not directly comparable to those of the above works; in particular, faint but broad components will have limited impact on the model line width and centroid velocity. On the other hand, single component fits can still robustly show spatial variation in kinematic properties, and a brightening of a broad component will manifest itself as a larger measured line width. Looking at the FWHM maps in the right panels of Figure 7, it is interesting to note that while we know from the stacked spectra that a broad component is present in the LCE cluster, it is not strong enough to lead to an elevation in the FWHM map; which is dominated by a component of km s-1 consistent with the surroundings.
@@
The line is however substantially wider in a small region immediately next to the LCE as seen in the two bright spikes in P1, reaching an FWHM of km s-1. The same region appears twice in P2+3 in highly elongated form due to the strong lensing shear in these images. This region is also slightly blueshifted relative to the LCE () and shows an equally conspicuous spike in the [N II]/H ratio. It is not immediately clear what drives this apparent outflow, as it is spatially distinct from the LCE and there is no star cluster visible at its location. It is close enough to the LCE to have bled significantly into the MagE and FIRE slits used by R. Mainali et al.. We therefore conclude from our maps that this region, rather than the LCE itself, gives rise to the broad component in the LCE spectra in that paper. The ground-based spectra of R. Mainali et al. (2022) lacked the spatial resolution to make this distinction. If correct, this interpretation of course raises the question of how often the observed broad components in LCEs from the literature originate from nearby regions, rather than from the isotropic stellar winds of the leaking clusters themselves, and in extension, whether a wider array of mechanisms are needed to explain the appearance and sustenance of ionizing escape channels.
@@ 5.2 Dust geometry
The leaking clump is known to have a bright, narrow Ly emission spike at systemic velocity (T. E. Rivera-Thorsen et al., 2017b). This is indicative of have a line of sight towards us that is virtually free of neutral gas. However, the dust reddening as traced by the Balmer decrement is not significantly lower along this line of sight compared to others in the vicinity. This lack of a hole in the dust, together with the relatively smooth and undramatic behavior of the kinematic parameters around the LCE, indicates the absence of strong mechanical feedback. If mechanical feedback is not at work, this implies that photoionization is the main mechanism creating the escape path for Ly and LyC photons, at least in the immediate surroundings of the LCE. As discussed above in Sect. 5.1, the suggests that the galaxy is currently or recently interacting with anothet galaxy. As discussed below, an interaction could have helped strip the bulk of the surrounding H i envelope away from the region surrounding the LCE.
@@ 5.3 Ionization
As discussed in section 1, the diagnostics is a good indicator of overall ionization in a given region, but not a good predictor of LyC escape. An alternative, ionization-based tracer of LyC escape is the relative weakness of [S II], first suggested by R. M. Alexandroff et al. (2015). Like the O32 diagnostic, it is a measure of ionization; but the ionization potential of neutral sulphur is as low as 10.36 eV, lower than the ionization potential of hydrogen, such that it is effectively shielded from ionization in neutral regions. [S II] lines are only emitted from only present at the outer boundary of H II regions. These boundaries disappear as a region becomes fully ionized, exactly in the conditions that allow LyC escape.
@@
For the Sunburst Arc, we can compare the spatially resolved maps of log [S II]/H shown in the rightmost panels of Figure 8 to both the local values at the LCE cluster, and the global, integrated value as found in the previous section. From Figure 8, it is clear to see in the rightmost panels that the [S II]/H map has a strong, and highly localized, minimum at all the images of the LCE cluster. From Table 1, we see that the local value at the LCE is indeed , consistent with what is seen in the figure, while the global value is . These latter two values correspond to values of the quantity [S II] as defined in B. Wang et al. (2021) of and , respectively. Comparing to the histograms and derived kernel density estimates in Fig. 2 in B. Wang et al. (2021), we see that these values put the LCE into a range in which the LzLCS contains practically no non-LCEs, while the integrated value falls in a tricky region, which on one hand is dominated by by non-leakers, but which on the other hand is also close to the peaks of both the distribution of strong LCEs and that of all LCEs in that sample.
Thus, the LCE cluster alone has such strong [S II] that it could not be mistaken for a non-leaker. In contrast, the galaxy as a whole displays a [S II] level typical for the majority of both leakers and non-leakers, and thus this quantity loses a great deal of its diagnostic strength when integrated over the whole galaxy, at least when compared to the LzLCS.
@@
The strong contrast in [S II] at the LCE compared to the surroundings also agrees very well with the predictions, based on which the method was initially proposed, and suggests that it is indeed well physically motivated.
@@
For comparison, Figure 10 also shows the values of [S II] from the stacked archival JWST spectra of star forming galaxie from M. J. Hayes et al. (2025). The stacks are grouped by EW([O III]) as a proxy of ionization. While the Sunburst Arc LCE stack sits somewhere between the fourth and fifth quintile of the high- ionization distribution, the global spectrum is more modest. Interestingly, though, even the galaxy as a whole has higher ionization and stronger sulfur deficit than the majority of observed galaxies in the high-{z}.
@@ 5.4 Global ionization properties
We also produced the classic Baldwin, Philips, Terlevich (BPT J. A. Baldwin et al., 1981; L. J. Kewley et al., 2001; G. Kauffmann et al., 2003; L. J. Kewley et al., 2006) diagram of the Sunburst Arc based on line ratios extracted both locally in the LCE cluster stack, and globally from the weighted map described in subsubsection 3.7.1, left panel. For comparison, we also included the LzLCS (B. Wang et al., 2021; S. R. Flury et al., 2022b), with their LCEs shown as filled and their non-LCEs as empty circles. The shaded zone in the upper left corner of the diagram is the D. K. Erb et al. (2016) Extreme Emission Line Galaxy (EELG) classification region. Interestingly, while the Sunburst Arc on the diagnostic maps in e.g. Figure 8 can look like an overall quite modestly ionized galaxy, once the diagnostics properties get weighted by flux as they would in a spatially integrated spectrum, it is revealed as a quite extreme galaxy, even for its redshift, which is noteworthy considering the general breakdown of the correlations between just these these properties and LyC at .
@@ 5.5 LyC escape scenario
T. E. Rivera-Thorsen et al. (2017b) observed a triple-peaked Ly profile emanating from the Sunburst Arc, consisting of the classic double-peak profile observed many times in the literature, combined with a narrow peak with a peak velocity and line width almost identical to those of H, telling of a component undergoing no radiative transfer effects, escaping through a channel of less then one optical depth in Ly. Based on this, T. E. Rivera-Thorsen et al. (2017b) predicted a scenario in which Ly escapes directly through a narrow perforation in an otherwise optically thick shroud of H i, in a manner reminiscent of the arc’s namesake meteorological phenomenon. According to our best models at the time, the total solid opening angle could be no more than than 2-5% of the total projected area surrounding the LCE cluster, with an extremely low neutral column density of . If the combined opening angle were larger, practically all Ly photons were predicted to escape through the open channel without the wavelength-space redistribution leading to the observed dual peaks surrounding the central peak. Such a channel would have to be created by a highly directional effect such as a jet from a previous accretion event; and it would have to be coincidentally aligned to a high degree with our line of sight. While such a scenario is certainly possible and might indeed have been the only plausible explanation; having to invoke this level of coincidental fine-tuning is uncomfortable, and furthermore, such a puncture by jet scenario would also have cleared away the majority of dust along the line of sight, but a significant dust screen was already then observed to be present in the nebular emission in T. E. Rivera-Thorsen et al. (2017b), and since confirmed in models of the stellar population (J. Chisholm et al., 2019; T. E. Rivera-Thorsen et al., 2019).
@@
Recent modeling work by S. Almada Monter & M. Gronke (2024) has shown that H i clouds are much more “sticky” to Ly photons than previously believed, in the sense that Ly photons have a substantially higher probability of getting trapped in the H i gas and undergoing significant spatial and frequency-space redistribution, than was believed at the time of T. E. Rivera-Thorsen et al. (2017b, 2019). Consequently, a much wider range of opening angle than previously believed can give rise to the characteristic Ly profile of the Sunburst Arc; perhaps angles as large as 30-40% of the total projected area. In this case, there is no longer a need to invoke a purely serendipetous, á priori very unlikely alignment of a jet or similar, highly dirctional, mechanism with our line of sight, and with this proposition also disappears the problem of explaining how such a case of mechanical feedback would have extremely efficiently cleared a pathway of neutral gas, yet left a seemingly unaltered dust screen behind. With the new anisotropy characteristics allowed by S. Almada Monter & M. Gronke (2024), we once again find ourselves in territory in which photoionization combined with an asymmetric H i geometry could clear away the H i in a cone-like configuration, while leaving the dust screen largely unchanged, as well as leaving behind a majority of projected surrounding area still containing enough H i to make it optically thick to Ly, thus generating the classic double-peaked Ly profile surrounding the central peak.
@@
We propose that this asymmetric H I geometry is provided by the same tidal stripping from a major interaction as found in Haro 11, discussed in section 1. Comparing Figure 2 to Fig. 1(a) in A. Le Reste et al. (2024), the Sunburst source galaxy has a geometry consistent with a similar scenario. In the upper (N) end is the interacting companion, connected to the main galaxy with a bridge-like structure in the central part of the figure. The LCE cluster is located in the bottom (S) end of the figure, farthest away from the companion in a region that would be stripped of much of its H i in a scenario similar to Haro 11. We therefore propose that tidal stripping, while not shown to actually take place, could plausibly be occurring also in the Sunburst galaxy and could well be at least in part responsible for clearing the paths for the escaping ionizing radiation. The high ionizing photon production rate in the LCE cluster has helped ionize the remaining neutral ISM with extreme efficiency, given the extremely low neutral column density inferred from the Ly profile. The seemingly even dust cover suggests that photoionization, rather than mechanical feedback or radiative pressure, has played the main role in clearing the ionizing escape path.
@@
A conceptual view of this escape scenario is shown in Figure 12. The figure shows in sketch form the same view as Figure 2. Here, the blue, green, and orange patches represent the main source galaxy, the faint bridge, and the redder interacting companion. The main stellar clumps are drawn in the same positions as seen in that figure. Superimposed is shown in grayscale an illustrative set of contours as they could plausibly look in a case of recent or ongoing tidal stripping by the companion (See A. Le Reste et al., 2024, Fig. 2, for a real-world observed scenario). The LCE cluster has ionized a surrounding H ii region to a degree where the surrounding neutral Hydrogen, already thin from the tidal interaction, has been completely ionized, allowing LyC photons to well out in a cone-like shape. This cone can, from the results of (S. Almada Monter & M. Gronke, 2024), be up to of the projected surrounding area and still be consistent with the Ly profile reported in (T. E. Rivera-Thorsen et al., 2017a).
@@ 6 Summary and conclusion
In this work, we present NIRSpec integral field spectroscopy of three pointings at the gravitationally lensed LyC-leaking starburst galaxy, the Sunburst Arc. For each spaxel, we have simultaneously fitted the strongest emission lines to each a single Gaussian profile, all with a shared redshift and FWHM in velocity space, yielding maps of velocity centroid, FWHM, and line fluxes at a pixel resolution of . We constructed maps dust reddening, [S II]
@@
Line centroid velocity maps show that the Sunburst Arc source galaxy displays a strong, though disturbed, velocity gradient, with a velocity span
@@
The kinematic maps show that the Sunburst Arc is not clearly dominated by neither rotation nor turbulence. The inclination-uncorrected shearing velocity is . If the inclination is more than 45°, which is the case for 70% of randomly distributed galaxies, it is dispersion-dominated, if not very strongly. In either case, the velocity dispersion is comparatively high for a star forming galaxy, indicating substantial perturbations. We interpret this as a sign of current and/or recent interaction with the companion seen in the upper part of Figure 2, which may have helped spark strong star formation in the main galaxy.
@@
In a stacked spectrum extracted from five gravitationally lensed images of the LCE cluster, we have identified more than 60 rest-frame optical and red emission lines, spanning from O III 3132 to Pa9 9229, including Balmer lines up to H20 and Paschen lines from Pa9 to Pa22. The stacked spectrum includes a Balmer jump, treated in (B. Welch et al., 2025).
@@
The LCE stack has a sulfur deficiency of . Compared to the LzLCS, this places the Sunburst Arc in a range almost exclusively populated by strong LyC emitters. In contrast, the integrated spectrum , a value which in the LzLCS is populated by both strong and weak leakers and non-leakers (B. Wang et al., 2021), showing that contributions from less ionized surrounding regions can dilute this signature even for a strong LyC leaker.
@@
The strong, broad line emission in [O III] and H observed by (R. Mainali et al., 2022) does most likely not originate in the LCE itself, but in a Nitrogen-enriched cloud in its vicinity, with a physical distance within a few tens of parsec. This could re-ignite the debates about the role of stellar winds and alternative momentum injection mechanisms in clearing the escape paths for LyC photons.
@@
Based on the extremely low H I column density(T. E. Rivera-Thorsen et al., 2017b), the presence of an interacting companion to the N of the main galaxy(K. Sharon et al., 2022), the strong spike in ionization around the LCE and lack of signs of strong mechanical feedback around the LCE cluster; we propose a LyC escape scenario in which tidal interaction with the companion has displaced the majority of the H I envelope around the LCE, leaving the rest to be easily ionized by the strong ionizing emission from the LCE cluster itself.

### v3

@@ 5.1.1 The origin of a broad emission component in H and [O III]
It has long been predicted that stellar winds and supernova feedback would play an important role in clearing the escape paths for Lyman-continuum, both in theoretical works (trebitsch2017; rosdahl2018; kakiichi2021) and in observations (heckman2011; chisholm2017; kim2020). amorin2024 have found a strong preponderance of broad emission lines in LyC-leaking Green Pea galaxies, although it has been difficult to establish any direct correlations between broad emission components and LyC escape fraction.
@@
mainali2022 studied the rest-UV and -optical emission in the Sunburst Arc from the ground using the Magellan/MagE and Magellan/FIRE spectrographs, respectively. These authors did find a strong broad component in the emission features of H and [O III] in the LCE spectra, while spectra from non-leaking regions showed a considerably fainter broad feature. However, the broad feature in mainali2022 has FWHM km s-1 or km s-1 somewhere in between the broad- and medium-width features of (amorin2024). Compared to the ground-based spatially unresolved spectra of mainali2022, the current NIRSpec IFU observations enable spectra to be extracted from spatial regions much more selectively, with lower contamination from nearby regions.
@@
riverathorsen2024; welch2025 extracted and stacked spectra from multiple lensed images of the LCE, allowing for unprecedented signal/noise ratio and absence of contamination from surrounding regions. Both these works modeled strong, rest-frame optical nebular emission lines using three Gaussian components, yielding FWHM values of km s-1; consistent with each other and broadly consistent with typical values of the narrow, medium and broad components found by amorin2024.
@@
In this work, we have only used one Gaussian component to model the emission lines in each spaxel, so our measured velocities are not directly comparable to those of the above works; in particular, faint but broad components will have limited impact on the model line width and centroid velocity. On the other hand, single component fits can still robustly show spatial variation in kinematic properties, and a brightening of a broad component will manifest itself as a larger measured line width. Looking at the FWHM maps in the right panels of Figure 7, it is interesting to note that while we know from the stacked spectra that a broad component is present in the LCE cluster, it is not strong enough to lead to an elevation in the FWHM map; which is dominated by a component of km s-1 consistent with the surroundings.
@@
The line is however substantially wider in a small region immediately next to the LCE as seen in the two bright spikes in P1, reaching an FWHM of km s-1. The same region appears twice in P2+3 in highly elongated form due to the strong lensing shear in these images. This region is also slightly blueshifted relative to the LCE () and shows an equally conspicuous spike in the [N II]/H ratio. It is not immediately clear what drives this apparent outflow, as it is spatially distinct from the LCE and there is no star cluster visible at its location. It is close enough to the LCE to have bled significantly into the MagE and FIRE slits used by mainali2022. We therefore conclude from our maps that this region, rather than the LCE itself, gives rise to the broad component in the LCE spectra in that paper. The ground-based spectra of mainali2022 lacked the spatial resolution to make this distinction. If correct, this interpretation of course raises the question of how often the observed broad components in LCEs from the literature originate from nearby regions, rather than from the isotropic stellar winds of the leaking clusters themselves, and in extension, whether a wider array of mechanisms are needed to explain the appearance and sustenance of ionizing escape channels.
@@ 5.2 Dust geometry
The leaking clump is known to have a bright, narrow Ly emission spike at systemic velocity (sunburst2017). This is indicative of have a line of sight towards us that is virtually free of neutral gas. However, the dust reddening as traced by the Balmer decrement is not significantly lower along this line of sight compared to others in the vicinity. This lack of a hole in the dust, together with the relatively smooth and undramatic behavior of the kinematic parameters around the LCE, indicates the absence of strong mechanical feedback. If mechanical feedback is not at work, this implies that photoionization is the main mechanism creating the escape path for Ly and LyC photons, at least in the immediate surroundings of the LCE. As discussed above in Sect. 5.1, the suggests that the galaxy is currently or recently interacting with anothet galaxy. As discussed below, an interaction could have helped strip the bulk of the surrounding H i envelope away from the region surrounding the LCE.
@@ 5.3 Ionization
As discussed in section 1, the diagnostics is a good indicator of overall ionization in a given region, but not a good predictor of LyC escape. An alternative, ionization-based tracer of LyC escape is the relative weakness of [S II], first suggested by alexandroff2015. Like the O32 diagnostic, it is a measure of ionization; but the ionization potential of neutral sulphur is as low as 10.36 eV, lower than the ionization potential of hydrogen, such that it is effectively shielded from ionization in neutral regions. [S II] lines are only emitted from only present at the outer boundary of H II regions. These boundaries disappear as a region becomes fully ionized, exactly in the conditions that allow LyC escape.
@@
For the Sunburst Arc, we can compare the spatially resolved maps of log [S II]/H shown in the rightmost panels of Figure 8 to both the local values at the LCE cluster, and the global, integrated value as found in the previous section. From Figure 8, it is clear to see in the rightmost panels that the [S II]/H map has a strong, and highly localized, minimum at all the images of the LCE cluster. From Table 1, we see that the local value at the LCE is indeed , consistent with what is seen in the figure, while the global value is . These latter two values correspond to values of the quantity [S II] as defined in wang2021 of and , respectively. Comparing to the histograms and derived kernel density estimates in Fig. 2 in wang2021, we see that these values put the LCE into a range in which the LzLCS contains practically no non-LCEs, while the value from the integrated spectrum places it slightly above the average from SDSS. This value falls in a tricky region in the KDEs of wang2021; it is on one hand dominated by by non-leakers, but on the other hand it is also close to the peaks of the distributions both of strong LCEs and that of all LCEs in the LzLCS.
Thus, the LCE cluster alone has such strong [S II] that it could not be mistaken for a non-leaker. In contrast, the galaxy as a whole displays a [S II] level which is almost equally represenative for leakers and non-leakers. Thus, we see that while [S II] seems to be a very good local tracer of ionized channels, the diagnostic can still be diluted by surrounding regions, making a leaker look like a non-leaker, as also found in wang2019; wang2021.
@@
The strong contrast in [S II] at the LCE compared to the surroundings also agrees very well with the predictions, based on which the method was initially proposed, and supports that the method is indeed well physically motivated.
@@
For comparison, Figure 10 also shows the values of [S II] from the stacked archival JWST spectra of star forming galaxie from hayes2025. The stacks are grouped by EW([O III]) as a proxy of ionization. While the Sunburst Arc LCE stack sits somewhere between the fourth and fifth quintile of the high- ionization distribution, the global spectrum is more modest. Interestingly, though, even the galaxy as a whole has higher ionization and stronger sulfur deficit than the majority of observed galaxies in the high-{z}.
@@ 5.4 Global ionization properties
We also produced the classic Baldwin, Philips, Terlevich (BPT baldwin1981; kewley2001; kauffmann2003; kewley2006) diagram of the Sunburst Arc based on line ratios extracted both locally in the LCE cluster stack, and globally from the weighted map described in subsubsection 3.7.1, left panel. For comparison, we also included the LzLCS (wang2021; flury2022b), with their LCEs shown as filled and their non-LCEs as empty circles. The shaded zone in the upper left corner of the diagram is the erb2016 Extreme Emission Line Galaxy (EELG) classification region. Interestingly, while the Sunburst Arc on the diagnostic maps in e.g. Figure 8 can look like an overall quite modestly ionized galaxy, once the diagnostics properties get weighted by flux as they would in a spatially integrated spectrum, it is revealed as a quite extreme galaxy, even for its redshift, which is noteworthy considering the general breakdown of the correlations between just these these properties and LyC at .
@@ 5.5 LyC escape scenario
sunburst2017 observed a triple-peaked Ly profile emanating from the Sunburst Arc, consisting of the classic double-peak profile observed many times in the literature, combined with a narrow peak with a peak velocity and line width almost identical to those of H, telling of a component undergoing no radiative transfer effects, escaping through a channel of less then one optical depth in Ly. Based on this, sunburst2017 predicted a scenario in which Ly escapes directly through a narrow perforation in an otherwise optically thick shroud of H i, in a manner reminiscent of the arc’s namesake meteorological phenomenon. According to our best models at the time, the total solid opening angle could be no more than than 2-5% of the total projected area surrounding the LCE cluster, with an extremely low neutral column density of . If the combined opening angle were larger, practically all Ly photons were predicted to escape through the open channel without the wavelength-space redistribution leading to the observed dual peaks surrounding the central peak. Such a channel would have to be created by a highly directional effect such as a jet from a previous accretion event; and it would have to be coincidentally aligned to a high degree with our line of sight. While such a scenario is certainly possible and might indeed have been the only plausible explanation; having to invoke this level of coincidental fine-tuning is uncomfortable, and furthermore, such a puncture by jet scenario would also have cleared away the majority of dust along the line of sight, but a significant dust screen was already then observed to be present in the nebular emission in sunburst2017, and since confirmed in models of the stellar population (chisholm2019; riverathorsen2019).
@@
Recent modeling work by almadamonter2024 has shown that H i clouds are much more “sticky” to Ly photons than previously believed, in the sense that Ly photons have a substantially higher probability of getting trapped in the H i gas and undergoing significant spatial and frequency-space redistribution, than was believed at the time of sunburst2017; riverathorsen2019. Consequently, a much wider range of opening angle than previously believed can give rise to the characteristic Ly profile of the Sunburst Arc; perhaps angles as large as 30-40% of the total surrounding solid angle. In this case, there is no longer a need to invoke a purely serendipetous, á priori very unlikely alignment of a jet or similar, highly directional, mechanism with our line of sight, and with this proposition also disappears the problem of explaining how such a case of mechanical feedback would have extremely efficiently cleared a pathway of neutral gas, yet left a seemingly unaltered dust screen behind. With the new anisotropy characteristics allowed by almadamonter2024, we once again find ourselves in territory in which photoionization combined with an asymmetric H i geometry could clear away the H i in a cone-like configuration, while leaving the dust screen largely unchanged, as well as leaving behind a majority of projected surrounding area still containing enough H i to make it optically thick to Ly, thus generating the classic double-peaked Ly profile surrounding the central peak.
@@
We propose that this asymmetric H I geometry is provided by the same kind of tidal stripping from a major interaction as found in Haro 11, discussed in section 1. Comparing Figure 2 to Fig. 1(a) in lereste2024, the Sunburst source galaxy has a geometry consistent with a similar scenario. In the upper (N) end is the interacting companion, connected to the main galaxy with a bridge-like structure in the central part of the figure. The LCE cluster is located in the bottom (S) end of the figure, farthest away from the companion in a region that would be stripped of much of its H i in a scenario similar to Haro 11. We therefore propose that tidal stripping, while not shown to actually take place, could plausibly be occurring also in the Sunburst galaxy and could well be at least in part responsible for clearing the paths for the escaping ionizing radiation. The high ionizing photon production rate in the LCE cluster has helped ionize the remaining neutral ISM with extreme efficiency, given the extremely low neutral column density inferred from the Ly profile. The seemingly even dust cover suggests that photoionization, rather than mechanical feedback or radiative pressure, has played the main role in clearing the ionizing escape path.
@@
A conceptual view of this escape scenario is shown in Figure 12. The figure shows in sketch form the same view as Figure 2. Here, the blue, green, and orange patches represent the main source galaxy, the faint bridge, and the redder interacting companion. The main stellar clumps are drawn in the same positions as seen in that figure. Superimposed is shown in grayscale an illustrative set of contours as they could plausibly look in a case of recent or ongoing tidal stripping by the companion (See lereste2024, Fig. 2, for a real-world observed scenario). The LCE cluster has ionized a surrounding H ii region to a degree where the surrounding neutral Hydrogen, already thin from the tidal interaction, has been completely ionized, allowing LyC photons to well out in a cone-like shape. This cone can, from the results of (almadamonter2024), be up to of the projected surrounding area and still be consistent with the Ly profile reported in (riverathorsen2017).
@@ 6 Summary and conclusion
In this work, we present NIRSpec integral field spectroscopy of three pointings at the gravitationally lensed LyC-leaking starburst galaxy, the Sunburst Arc. For each spaxel, we have simultaneously fitted the strongest emission lines to each a single Gaussian profile, all with a shared redshift and FWHM in velocity space, yielding maps of velocity centroid, FWHM, and line fluxes at a pixel resolution of . We constructed maps of dust reddening, ionization parameter , ionization tracer [S II], and shock- and abundance tracer [N II]/H.
@@
Line centroid velocity maps show that the Sunburst Arc source galaxy displays a strong, though disturbed, velocity gradient, with an inclination- and lensing-uncorrected velocity span , and a lensing-uncorrected intrinsic velocity dispersion .
@@
The kinematic maps show that the Sunburst Arc is not clearly dominated by neither rotation nor turbulence. The ratio of shearing velocity to intrinsic velocity dispersion is . If the inclination is more than 45°, which is the case for 70% of randomly distributed galaxies, it is dispersion-dominated. In either case, the velocity dispersion is comparatively high for a star forming galaxy, indicating substantial perturbations. We interpret this as a sign of current and/or recent interaction with the companion seen in the upper part of Figure 2, which may have helped spark strong star formation in the main galaxy.
@@
In a stacked spectrum extracted from five gravitationally lensed images of the LCE cluster, we have identified more than 60 rest-frame optical and red emission lines, spanning from O III 3132 to Pa9 9229, including Balmer lines up to H20 and Paschen lines from Pa9 to Pa22. The stacked spectrum includes a Balmer jump, treated in (welch2025).
@@
The LCE stack has a sulfur deficiency of . Compared to the LzLCS, this places the Sunburst Arc in a range almost exclusively populated by strong LyC emitters. In contrast, the integrated spectrum , a value slightly above the average, in a range which in the LzLCS is populated by both strong and weak leakers and non-leakers (wang2021), showing that contributions from less ionized surrounding regions can dilute this signature beyond detectability even for a strong LyC leaker.
@@
The strong, broad line emission in [O III] and H observed by (mainali2022) does most likely not originate in the LCE itself, but in a Nitrogen-enriched cloud in its vicinity, with a physical distance within a few tens of parsec. This could re-ignite the debates about the role of stellar winds and alternative momentum injection mechanisms in clearing the escape paths for LyC photons.
@@
Based on the extremely low H I column density along the LOS to the LCE clump (sunburst2017), the presence of an interacting companion to the N of the main galaxy(sharon2022), the strong spike in ionization around the LCE and lack of signs of strong mechanical feedback around the LCE cluster; we propose a LyC escape scenario in which tidal interaction with the companion has displaced the majority of the H I envelope around the LCE, leaving the rest to be easily ionized by the strong ionizing emission from the LCE cluster itself.